Robbie Roberts, owner of CupAMoes Coffee, is trying to determine where to locate his newest distribution warehouse. Robbie wCapital city (Pop. 400,000) (1.0.5.0) Springfield (Pop. 200,000) (4.5, 3.0) Shelbyville (Pop. 170,000) (4.0, 1.0) 4
Answer #1

Answers @ Total population Wa 400000 + 200000+ 110000 - 770000 Total xw = (400000X1)+(200000 x 4.5 )+(170000 x 4) = 400000 902770000 770 000 = 3.60 ogue = hos oix coordinale = 2.57 & y coordinate = 3,6 | Total sales w=800000 + 15000ot 15oooo 1100 0004600000 1100000 (y = 4.18 Distribution centre 1xcoordinate x ly coordinatiy salesw. 1 xw YW 800000 800000 4000000 capital cit